# Simultaneous global and local clustering in multiplex networks with covariate information

### v1

@@ 3.1 Variational approximation of the posterior distribution for the HMPSBM
×$\times\prod_{k=1}^{M_{w}}q(\bm{\varphi}_{k}^{0})\times\prod_{k=1}^{M_{w}}q(\sigma_{k}^{2})\times\prod_{k=1}^{M_{w}}\prod_{s=1}^{M_{z}}q(\gamma_{ks}^{\prime}).$ (3.8)
A variational approximation is posited directly upon the stick-breaking variables of the GEM distribution, following the approach of Blei and Jordan (2006), and two truncation parameters $M_{z},M_{w}\in\mathbb{N}$ are introduced, implying that $q(\rho_{km})=\delta_{0}(\rho_{km})$ for $\{k,m\in\mathbb{N}:\max\{k,m\}>M_{z}\}$, $q(\bm{\varphi}_{k})=\delta_{0}(\bm{\varphi}_{k})$ for $\{k\in\mathbb{N}:k>M_{w}\}$, $q(\bm{\varphi}_{k}^{0})=\delta_{0}(\bm{\varphi}_{k}^{0})$ for $\{k\in\mathbb{N}:k>M_{w}\}$, $q(\sigma_{k}^{2})=\delta_{0}(\sigma_{k}^{2})$ for $\{k\in\mathbb{N}:k>M_{w}\}$, and $q(\gamma_{ks}^{\prime})=\delta_{0}(\gamma_{ks}^{\prime})$ for $\{k,s:k>M_{w}\mid m>M_{z}\}$. Here we define $\delta_{0}(\cdot)$ to be the Dirac delta function at 0. Note that the model set-up of (2.6)-(2.14) is infinite-dimensional, and we only make a finite approximation to it in the variational procedure.
@@
[…]
@@
[…]
@@
[…]
@@ 3.2 Initialisation
The parameters $\tilde{\alpha}_{\gamma_{ks}}$ and $\tilde{\beta}_{\gamma_{ks}}$ are initialised by computing the proportion of nodes $i\in\mathcal{V}$ that have $w_{i}=k$ and $z_{\ell i}=s$ This value is set equal to $\tilde{\alpha}_{\gamma_{ks}}/(\tilde{\alpha}_{\gamma_{ks}}+\tilde{\beta}_{\gamma_{ks}})$ and inverted to solve for $\tilde{\alpha}_{\gamma_{ks}}$, where we again first set $\tilde{\beta}_{\gamma_{ks}}=1$.
@@ 4 Simulation studies
We evaluate the performance of the inference procedure proposed in Section 3 using simulated data. In Section 4.1, we examine recovery of the global and layer-level groups with different truncation parameters. Section 4.2, considers a network with well-separated layer-level behaviour but with increasingly uninformative nodal covariates. Section 4.3 considers networks with informative features, but decreasing separation between each global group’s behaviour across layers. Next, Section 4.4 examines group recovery with varying network size and varying numbers of layers, in particular considering a setting in which our connectivity matrix has weak separation. We evaluate using the normalised mutual information (NMI; Strehl and Ghosh, 2002) score, which ranges from 0 to 1, with 0 corresponding to random assignment, and 1 perfect agreement. This metric accounts for labelling differences between simulations.
@@ 4.1 Inference of the number of latent groups
and run the CAVI inference procedure for 10 iterations. We run the procedure 50 times with $M_{w}=2$ and $M_{z}=3$, and again with $M_{w}=M_{z}=5$. In Table 2 we see that the NMI for the global groups is very high for both truncation parameter pairs, with a median value of 1. A corresponding table for the layer level groups is not included as the median was found to be 1.0 with a standard deviation of 0, indicating perfect recovery in all simulations.
@@ 4.2 Decreasing feature separation
where $\bm{\mu}_{k}$ corresponds to global group $k$. We allow $\alpha$ to vary over the set $\{5/2,2,3/2,1,1/2,0\}$, corresponding to decreasing separation with decreasing $\alpha$. The connection probability matrix $\rho$ we set to be as in the simulation of Section 4.1. The CAVI scheme was run for 15 iterations, which was found to be sufficient for convergence in the sense that relative changes in consecutive ELBO evaluations fell below a threshold $\epsilon$.
@@ 4.4 Effect of the number of layers and network size
Figure 4: Boxplots of the NMI score with varying number of nodes, $N$, and varying numbe of samples, $M$, for the global and layer groups under the settings of Section 4.4.
@@ 5 Application to the FAO trade network
The network does not contain nodal covariates, and so we source features from the World Bank (World Bank, 2023). We include the GDP from 2009 of each country and its urban land area in square kilometers, and we use a log-transformation before normalising to have zero mean and unit variance. We also include an intercept term for each node. For some countries, such as Macao and the Netherlands Antilles (which was dissolved in 2010 into separate Caribbean territories), we were able to get only one feature. We sorted each feature, found the rank of the feature we could obtain and used that ranking to impute the missing value of the other feature. One country is namedUnspecified, and for this we impute the values of the covariates using the global median. The algorithm of Section 3 was run with varying truncation parameters, and we found that the number of layer-level groups consistently collapsed onto 6. We thus set $M_{w}=12$ and $M_{z}=6$ in the inference and allow for 25 iterations of the CAVI procedure with a maximum of 30 gradient ascent steps per parameter. The algorithm detected 11 global groups and 6 layer-level groups (the same numbers were found with higher truncation values).
@@
In Figure 5 we see the distribution of layer-level groups within each global group. This corresponds to an estimate of $\{\bm{\gamma}_{k}\}_{k=1}^{11}$ in the model. We see that the behaviour of each global group is distinct from one another, with different global groups having different distributions across the layer-level groups.
@@
Figure 5: Distribution of layer-level groups in each global group.
@@
In Figure 6 is a world map with countries that were present in the network coloured by their inferred global group. Group 4 (dark green) consists of France, China (main land), the United States of America, the United Kingdom, Italy, Spain, the Netherlands and Germany. This is an intuitive community as they represent some of the largest global economies and trading powers at the time. Group 3 (light green) consists of many emerging economies and regional powers (Brazil, Russia, Argentina) plus some smaller EU economies (Portugal, Greece) and Asian trading centers (Hong Kong, Malaysia). Group 2 seems to sit between these two groups, consisting of a mixture of advanced economies (Japan, Switzerland, Denmark, Sweden, Canada, South Korea) and emerging economies with major agricultural output (India, Turkey, Mexico, South Africa). A particularly interesting community is the group 8 (dark orange) which is made-up of only Syria and Iran. The dataset in question is from the year before the Syrian Revolution of 2011. Iran was one of the largest financial supporters of then President Bashar al-Assad and so it is interesting to see that our inference procedure has clustered these countries together.
@@
We then use t-Distributed Stochastic Neighbor Embedding (t-SNE), a popular non-linear dimension reduction tool, to visualise our layer-level clustering. Figure 7 shows plots of a two-dimensional t-SNE embedding for layers 1 to 8 of the network, with nodes coloured according to their layer-level group. We see that nodes of the same colour tend to appear near to one another the embedding space, and that this holds in all of the layers of the network. This demonstrates that the clustering has effectively uncovered latent structure in the multiplex network. An analogous plot for all layers of the network is provided in Figure 8 in Appendix E.
Most of the panels in Figure 7 (with the exception of Layer 8) show a conic shape, with the blue groups (layer groups 1 and 2) showing low variance, the reds (layer groups 5 and 6) high variance and the greens (layer groups 3 and 4) sitting between the two. In Figure 5, we see that global groups 2 and 4 are mostly concentrated on layer groups 1 and 2. These global groups correspond mostly to advanced economies, which explains their low variability. Similarly, global groups 1, 9 and 11 are emerging economies in Africa and are concentrated on layer groups groups 5 and 6, which corresponds to the higher variance of the clusters.
@@ A.1 Complete conditional distribution for $z_{\ell i}$
×$\times\rho_{mk}^{A_{\ell ji}}(1-\rho_{mk})^{1-A_{\ell ji}}\bigg{]}^{\mathbb{I}\{z_{\ell i}=k,z_{\ell j}=m\}},$ (A.5)
@@
×$\times\rho_{mk}^{A_{\ell ji}}(1-\rho_{mk})^{1-A_{\ell ji}}\bigg{]}^{\mathbb{I}\{z_{\ell j}=m\}}\Bigg{\}}^{\mathbb{I}\{z_{\ell i}=k\}}$ (A.7)

### v2

@@ 3.1 Variational approximation of the posterior distribution for the HMPSBM
$\times\prod_{k=1}^{M_{w}}q(\bm{\varphi}_{k}^{0})\times\prod_{k=1}^{M_{w}}q(\sigma_{k}^{2})\times\prod_{k=1}^{M_{w}}\prod_{s=1}^{M_{z}}q(\gamma_{ks}^{\prime}).$ (3.8)
A variational approximation is posited directly upon the stick-breaking variables of the GEM distribution, following the approach of Blei and Jordan (2006), and two truncation parameters $M_{z},M_{w}\in\mathbb{N}$ are introduced, implying that $q(\rho_{km})=\delta_{0}(\rho_{km})$ for $\{k,m\in\mathbb{N}:\max\{k,m\}>M_{z}\}$, $q(\bm{\varphi}_{k})=\delta_{0}(\bm{\varphi}_{k})$ for $\{k\in\mathbb{N}:k>M_{w}\}$, $q(\bm{\varphi}_{k}^{0})=\delta_{0}(\bm{\varphi}_{k}^{0})$ for $\{k\in\mathbb{N}:k>M_{w}\}$, $q(\sigma_{k}^{2})=\delta_{0}(\sigma_{k}^{2})$ for $\{k\in\mathbb{N}:k>M_{w}\}$, and $q(\gamma_{ks}^{\prime})=\delta_{0}(\gamma_{ks}^{\prime})$ for $\{k,s\in\mathbb{N}:k>M_{w}\mid m>M_{z}\}$. Here we define $\delta_{0}(\cdot)$ to be the Dirac delta function at 0. Note that the model set-up of (2.6)-(2.14) is infinite-dimensional, and we only make a finite approximation to it in the variational procedure.
@@
    $\tilde{\bm{\phi}}_{w_{i},w}\propto\exp\{$ $\sum_{\ell\in[L]}\sum_{k=1}^{M_{z}}\bm{\phi}_{z_{\ell i},k}\left(\psi(\tilde{\alpha}_{\gamma_{wk}})-\psi(\tilde{\alpha}_{\gamma_{wk}}+\tilde{\beta}_{\gamma_{wk}})+\sum_{s=1}^{k-1}\left[\psi(\tilde{\beta}_{\gamma_{ws}})-\psi(\tilde{\alpha}_{\gamma_{ws}}+\tilde{\beta}_{\gamma_{ws}})\right]\right)$ (3.9)
@@
    $\tilde{\bm{\phi}}_{z_{\ell i},k}\propto\exp\{$ $\sum_{w=1}^{M_{w}}\tilde{\bm{\phi}}_{w_{i},w}\left(\psi(\tilde{\alpha}_{\gamma_{wk}})-\psi(\tilde{\alpha}_{\gamma_{wk}}+\tilde{\beta}_{\gamma_{wk}})+\sum_{s=1}^{k-1}\left[\psi(\tilde{\beta}_{\gamma_{ws}})-\psi(\tilde{\alpha}_{\gamma_{ws}}+\tilde{\beta}_{\gamma_{ws}})\right]\right)$ (3.11)
@@
where we write $\lfloor\cdot\rfloor$ to denote the strictly lower-diagonal part of a matrix. The log-Cholesky is preferred over a standard Cholesky parametrisation due to its superior numerical stability. Derivatives of the ELBO with respect to $B_{k}$ then become
@@
It should be noted that the $\bm{\varphi}_{k}$ will not be recoverable. The map $\alpha\mapsto\Phi(\alpha)$ is numerically non-invertible in the sense that it is not in practice a one-to-one function. For example, we have $\Phi(5)$ and $\Phi(6)$ agree to six decimal places. As the inference only ever processes the image of $\bm{x}_{i}^{\top}\bm{\varphi}_{k}$ under this mapping, we cannot hope to recover $\bm{\varphi}_{k}$. Furthermore, it is implicitly assumed that the features exhibit a group structure. In the case that the number of these groups is less than $P$, the design matrix $X$ will have an effective rank that is non-maximal. The result of this is that the map $\bm{\varphi}_{k}\mapsto X\bm{\varphi}_{k}$ has non-trivial nullspace, and so we cannot hope to recover $\bm{\varphi}_{k}$ uniquely. Furthermore, although the stick–breaking process imposes an order on the breaks, it does not fix a meaningful scale on the parameters as they are re-normalised in the inference. This, however, is not problematic for the inference as the primary object is the group structure of the nodes. In Section 4, we observe that the covariates do aid inference, despite being unable to recover $\bm{\varphi}_{k}$.
@@ 3.2 Initialisation
The parameters $\tilde{\alpha}_{\gamma_{ks}}$ and $\tilde{\beta}_{\gamma_{ks}}$ are initialised by computing the proportion of nodes $i\in\mathcal{V}$ that have $w_{i}=k$ and $z_{\ell i}=s$ . This value is set equal to $\tilde{\alpha}_{\gamma_{ks}}/(\tilde{\alpha}_{\gamma_{ks}}+\tilde{\beta}_{\gamma_{ks}})$ and inverted to solve for $\tilde{\alpha}_{\gamma_{ks}}$, where we again first set $\tilde{\beta}_{\gamma_{ks}}=1$.
@@ 4 Simulation studies
We evaluate the performance of the inference procedure proposed in Section 3 using simulated data. In Section 4.1, we examine recovery of the global and layer-level groups with different truncation parameters. Section 4.2, considers a network with well-separated layer-level behaviour but with increasingly uninformative nodal covariates. Section 4.3 considers networks with informative features, but decreasing separation between each global group’s behaviour across layers. Next, Section 4.4 examines group recovery with varying network size and varying numbers of layers, in particular considering a setting in which our connectivity matrix has weak separation. We evaluate using the normalised mutual information (NMI; Strehl and Ghosh, 2002) score, which ranges from 0 to 1, with 0 corresponding to random assignment, and 1 perfect agreement. This metric accounts for labelling differences between simulations. To align the inferred groups with the true simulated groups, we use an implementation of a modified Jonker-Volgenant algorithm with no initialisation (Crouse, 2016), using the linear_sum_assignment method from scikit-learn to re-enumerate the learned labels so that they are consistent with those of the simulation. The number of components are selected by assigning each node to the group corresponding to the index of its posterior probability vector with maximum value.
@@ 4.1 Inference of the number of latent groups
and run the CAVI inference procedure for 10 iterations, which was found to be sufficient for convergence in the sense that relative changes in consecutive ELBO evaluations fell below a small threshold $\epsilon$. The fast convergence time is attributed to the informed initialisation scheme we implement. We run the procedure 50 times with $M_{w}=2$ and $M_{z}=3$, and again with $M_{w}=M_{z}=5$. In Table 2 we see that the NMI for the global groups is very high for both truncation parameter pairs, with a median value of 1. A corresponding table for the layer level groups is not included as the median was found to be 1.0 with a standard deviation of 0, indicating perfect recovery in all simulations.
@@ 4.2 Decreasing feature separation
where $\bm{\mu}_{k}$ corresponds to global group $k$. We allow $\alpha$ to vary over the set $\{5/2,2,3/2,1,1/2,0\}$, corresponding to decreasing separation with decreasing $\alpha$. The connection probability matrix $\rho$ we set to be as in the simulation of Section 4.1.
@@ 4.4 Effect of the number of layers and network size
Figure 4: Boxplots of the NMI score with varying number of nodes, NN, and varying numbe of samples, $M$, for the global and layer groups under the settings of Section $4.$
@@ 5 Application to the FAO trade network
The network does not contain nodal covariates, and so we source features from the World Bank (World Bank, 2023). We run the inference twice, once using the GDP from 2009 of each country and its rural land area in square kilometers, and another using these two features, but additionally urban land area and coast-line length. We use a log-transformation before normalising to have zero mean and unit variance. We refer to these settings as setting A and setting B, respectively. We focus primarily on setting A, but will consider the results of setting B for comparison. We also include an intercept term for each node. For some countries, such as Macao and the Netherlands Antilles (which was dissolved in 2010 into separate Caribbean territories), we were able to get only one feature. We sorted each feature, found the rank of the feature we could obtain and used that ranking to impute the missing value of the other feature. One country is named Unspecified, and for this we impute the values of the covariates using the global median. The algorithm of Section 3 was run with varying truncation parameters, and we found that the number of layer-level groups consistently collapsed onto 6. We thus set $M_{w}=12$ and $M_{z}=6$ in the inference and allow for 25 iterations of the CAVI procedure with a maximum of 30 gradient ascent steps per parameter. The algorithm detected 11 global groups and 6 layer-level groups (the same numbers were found with higher truncation values).
@@
In Figure 5(a) we see the distribution of layer-level groups within each global group for setting A. This corresponds to an estimate of $\{\bm{\gamma}_{k}\}_{k=1}^{11}$ in the model. We see that the behaviour of each global group is distinct from one another, with different global groups having different distributions across the layer-level groups.
(a) Rural land area and GDP.
(b) Rural land area, GDP, urban land area and coast-line length.
@@
Figure 5: Distribution of layer-level groups in each global group using different features.
@@
In Figure 6 is a world map with countries that were present in the network coloured by their inferred global group for setting A. Group 4 (dark green) consists of France, China (main land), the United States of America, the United Kingdom, Italy, Spain, the Netherlands and Germany. This is an intuitive community as they represent some of the largest global economies and trading powers at the time. Group 3 (light green) consists of many emerging economies and regional powers (Brazil, Russia, Argentina) plus some smaller EU economies (Portugal, Greece) and Asian trading centers (Hong Kong, Malaysia). Group 2 seems to sit between these two groups, consisting of a mixture of advanced economies (Japan, Switzerland, Denmark, Sweden, Canada, South Korea) and emerging economies with major agricultural output (India, Turkey, Mexico, South Africa). A particularly interesting community is the group 8 (dark orange) which is made-up of only Syria and Iran. The dataset in question is from the year before the Syrian Revolution of 2011. Iran was one of the largest financial supporters of then President Bashar al-Assad and so it is interesting to see that our inference procedure has clustered these countries together.
@@
We then use t-Distributed Stochastic Neighbor Embedding (t-SNE), a popular non-linear dimension reduction tool, to visualise our layer-level clustering for setting A. Figure 7 shows plots of a two-dimensional t-SNE embedding for layers 1 to 8 of the network using a perplexity of 30, with nodes coloured according to their layer-level group. We see that nodes of the same colour tend to appear near to one another the embedding space, and that this holds in all of the layers of the network. This demonstrates that the clustering has effectively uncovered latent structure in the multiplex network. An analogous plot for all layers of the network is provided in Figure 8 in Appendix E.
Most of the panels in Figure 7 (with the exception of Layer 8) show a conic shape, with the blue groups (layer groups 1 and 2) showing low variance, the reds (layer groups 5 and 6) high variance and the greens (layer groups 3 and 4) sitting between the two. In Figure 5(a), we see that global groups 2 and 4 are mostly concentrated on layer groups 1 and 2. These global groups correspond mostly to advanced economies, which explains their low variability. Similarly, global groups 1, 9 and 11 are emerging economies in Africa and are concentrated on layer groups groups 5 and 6, which corresponds to the higher variance of the clusters.
We now examine the output for setting B. After aligning labels, the global-group NMI is 0.89 when compared to setting A, indicating very high agreement between the two fits. Figures 5(a) and 5(b) show that the layer group distributions within each global group are nearly unchanged between settings A and B, with the main difference in Group 10.
In total, 29 of the 177 countries (16.4%) changed global group after alignment, with 17 moving from Group 10 to Group 9. The countries driving this shift are mostly coastal/island states, consistent with the addition of coastline and urbanisation features. Four other countries moved into group 9 (Albania, Bolivia, Jamaica, Venezuela), three of which are coastal/island, again consistent with coastline length drawing them into a coastal cluster. We see stability within the developed-economy blocks. Importantly, the group corresponding to larger, more developed economies is unchanged by the added covariates: United States, United Kingdom, Germany, France, Italy, Spain, China, and the Netherlands all remain in the same global group as before. Likewise, other major economies such as Canada, Australia, Japan, Denmark, Sweden, and Switzerland also retain their original assignments. This indicates that the covariate augmentation refines the periphery (e.g., coastal/island states) while leaving the core, high-income structure intact that is predominantly determined by the network structure. Overall, the resulting change from setting A to setting B suggest that network structure remains the primary driver of the clustering, with the new covariates making small and interpretable refinements.
@@ A.1 Complete conditional distribution for $z_{\ell i}$
$\times\rho_{mk}^{A_{\ell ji}}(1-\rho_{mk})^{1-A_{\ell ji}}\bigg]^{\mathbb{I}\{z_{\ell i}=k,z_{\ell j}=m\}},$ (A.5)
@@
$\times\rho_{mk}^{A_{\ell ji}}(1-\rho_{mk})^{1-A_{\ell ji}}\bigg]^{\mathbb{I}\{z_{\ell j}=m\}}\Bigg\}^{\mathbb{I}\{z_{\ell i}=k\}}$ (A.7)
@@ Appendix D Full algorithm of CAVI with Adam
1 CAVI with Adam. Write $f_{k}(\bm{\theta},\Sigma)=\partial\mathcal{L}/\partial\tilde{\bm{\theta}}_{\varphi_{k}}$ evaluated at $\tilde{\bm{\theta}}_{\varphi_{k}}=\bm{\theta}$ and $\widetilde{{\Sigma}}_{\varphi_{k}}=\Sigma$ conditional upon all other current parameter estimates, and similarly $g_{k}(\bm{\theta},\Sigma)=\partial\mathcal{L}/\partial\widetilde{\Sigma}_{\varphi_{k}}$, as in (3.17)-(3.19). Write $\mathcal{L}(\bm{\theta},\Sigma)$ for the ELBO evaluated at $\tilde{\bm{\theta}}_{\varphi_{k}}=\bm{\theta}$ and $\widetilde{\Sigma}_{\varphi_{k}}=\Sigma$.